OpenStudy (anonymous):

Giving Medal For Answer...help? The surface area of the top surface of the water in a circular swimming pool is about 206 square feet. Estimate the radius of the pool, to the nearest foot. a. about 14 feet b. about 8 feet c. about 11 feet d. about 10 feet

OpenStudy (anonymous):

Area of circle is Pi * r * r so 206 = pi * r^2 so radius = \sqrt{206/pi} so it about 8 feet

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw